# NSD-DIL: Null-Shot Deblurring Using Deep Identity Learning

### v1

@@ Abstract
In this paper, we propose to reformulate the blind image deblurring task to directly learn an inverse of the degradation model using a deep linear network. We introduce “Deep Identity Learning (DIL)”, a novel learning strategy that includes a dedicated regularization term based on the properties of linear systems, to exploit the identity relation between the degradation and inverse degradation models. The salient aspect of our proposed framework is it neither relies on a deblurring dataset nor a single input blurred image (like Polyblur, a self-supervised method). Since it is purely image-data-independent, we term our model as “Null-Shot deblurring Using Deep Identity Learning (NSD-DIL)". We also provide an explicit representation of the learned deep linear network in a matrix form, called “Deep Restoration Kernel (DRK)" for deblurring task. The proposed framework detours the typical degradation kernel estimation step involved in most of the existing blind deblurring solutions by the proposition of our Random Kernel Gallery (RKG) dataset. In this work, we focus on the restoration of mild blur images, generated by small out-of-focus, lens blur, or slight camera motion, which often occurs in real images. Our experiments show that the proposed method outperforms both traditional and deep learning based deblurring methods, with at least an order of 100 lesser computational resources. The proposed NSD-DIL method can be effortlessly extended to the Image Super-Resolution (ISR) task as well to restore the low-resolution images with fine details. The NSD-DIL model and its kernel form representation (DRK) are lightweight yet robust and restore the mild blur input in a fraction of a second. Hence, more suitable for wide real-time applications.
@@
Convolution Deblurring Deconvolution Degradation kernel
@@ 1 Introduction
Image deblurring is a well-established low-level vision task whose objective is to generate a sharp image from the given corresponding blurred input observation(s). Blur can be caused by various factors like, but not limited to, camera shake, observed relative motion between camera and target, camera defocus, camera lens aberration, light diffraction due to finite lens aperture, integration of light in the sensor [10]. Deblurring is a severely ill-posed problem as multiple combinations of a sharp image and a blur kernel can result in the same blurred image. However, real-world applications in various prominent domains like medical imaging, satellite imaging, and surveillance demand a sharp version of the scene of interest for its analysis.
@@
Image deblurring solutions can be categorized into non-blind and blind approaches. Non-blind methods aim to restore clear images from blurred inputs when the degradation kernel is known. In contrast, blind deblurring methods work with an unknown degradation kernel. The traditional non-blind and blind deblurring methods include like Wiener filter [55], Richardson-Lucy algorithm [43], sophisticated priors [54, 26, 33], Maximum a-Posterior (MAP) approaches with a variety of data terms [42, 11] and prior terms [23, 48, 57, 65]. However, the optimization problems of these sophisticated model limit their practical usage. The blind deblurring works either estimate both the blur kernel and the sharp image simultaneously [33, 24, 4, 38] or the blur kernel is estimated first and then non-blind deblurring is performed [14, 8, 35]. It is demonstrated that estimating the blur kernel is easier than jointly estimating the blur kernel and sharp image together [27]. With the advent and continuous advancements of Deep Learning (DL) techniques, remarkable improvement in the restored sharp image quality has been observed. The recent deblurring solutions, including the pioneering works [50, 34], adopt sophisticated techniques like Residual blocks [40, 51, 61], Generative adversarial networks [25, 62], Recurrent Neural Networks [60, 37], Transformers [30, 59], and specialized loss functions [46, 63].
The major limitation, in non-blind methods is the error-free assumption on the blur kernel. In practice, the estimated blur kernel is erroneous to some extent which could cause severe artifacts, and degradations in the generated images [17, 29]. In the blind deblurring methods, the extensive dependency on datasets for prior information could limit their generalization ability. Besides these drawbacks, the present State-of-the-Art (SotA) image deblurring methods employ DL techniques using a large amount of image data with heavy computational and time complexities. Hence, not suitable for real-time applications [10, 7].
In recent literature, it is observed that most of the works were leaning towards addressing the strong motion and non-uniform blur [29, 41, 25, 18, 56]. However, very little attention was paid to the most common and practical use case where the blur is small i.e. mild blur that is often present in real images which degrades the image quality. Mild blur can be harmful to the performance of other image processing tasks, such as super-resolution. These tasks often rely on sharp images to produce accurate results. The works [10, 64], detached from the mainstream deblurring methods and were proposed solutions towards addressing the mild blur. For instance, Polyblur [10] is a blind and self-supervised approach that uses polynomial functions to restore image details blurred by mild camera shake or motion blur supporting real-time applications. It estimates the blur kernel of the image by analyzing the distribution of the gradient in the image and then applies the kernel multiple times, based on the severity of the blur.
@@
In this paper, we redefine the task of learning the deblurring model from image data to merely computing an inverse of the degradation model. We characterized the degradation model with a wide set of anisotropic Gaussian kernels perturbed with noise, called the Random Kernel Gallery (RKG) dataset. Here, we propose a novel learning strategy called “Deep Identity Learning (DIL)" to compute the inverse of the constructed degradation model i.e. RKG dataset. DIL exploits the identity relation between the degradation model and its inverse i.e. $K^{-1}$, deblur model. In this work, a custom lightweight and computationally efficient CNN with no activation functions, referred to as Linear-CNN (L-CNN), is trained on the RKG dataset with DIL as its objective. DIL includes a regularization term based on the frequency and spatial domain properties of the convolution operator to effectively ensure the identity system properties of our inverse degradation model. Unlike existing blind self-supervised deblur frameworks like [10], the proposed deblurring framework is image-independent and doesn’t require blurred or pair of blurred and sharp image data at any stage of the deblur task learning. Hence, we term the proposed model as “Null-Shot image Deblurring using Deep Identity Learning (NSD-DIL)". In this work, we provide a compact and effective representation of the proposed L-CNN, in a matrix form, called ”Deep Restoration Kernel (DRK)". To perform the deblur task, the L-CNN can be replaced with the DRK at the inference stage. This step transforms the deblur task into a simple convolution operation between the given blurred input image and the DRK. Therefore, the deblur task with the DRK further reduces the computational complexity in terms of time and memory by several orders. The proposed DRK is the convolution inverse of the degradation model and performs deblur tasks across a wide range of degradations more effectively. The highlights of our proposition are as follows.
@@
    To the best of our knowledge, NSD-DIL is the first DL-based image data-independent deblurring framework.
@@
    We propose a novel learning function i.e. DIL based on identity system properties with a dedicated regularization term driven by the spatial and frequency domain inferences to effectively learn an inverse degradation model.
@@
  4. 4.
    The proposed NSD-DIL model and the DRK perform superior deblurring, in case of mild blur, against the existing SotA deblur methods while being at least 100 times more computationally efficient.
@@
    We do demonstrate that the proposed NSD-DIL framework can be readily extended to the Image Super-Resolution (ISR) task.
The overall organization of the rest of the paper is as follows. Section 2 presents Related work. A detailed explanation of our proposed NSD-DIL method is given in Section 3. In Section 4, a thorough experimental setup and results are presented, followed by a Discussion in Section 5 and a Conclusion in Section 6.
@@ 2 Related work
Besides the blind deblurring works mentioned in the Sec 1, the other related works are as follows.
Image sharpening methods are simple and intuitive to generate sharp images. These methods share a common framework, the input image is filtered and then a proportion of the residual image is added back to the original input. This procedure boosts high-frequencies and increases the contrast in the generated image. The most well-established works include Unsharp Masking and its adaptations. These are simple and effective but sensitive to noise and generate artifacts [39, 45, 21]. Bilateral filter [53] and its variants [15, 3] were used to enhance the local contrast and high-frequency details in an image. Based on local structure and, sharpness information [64] proposed an adaptive sharpening method that performs simultaneous noise reduction and sharpening. However, the deblurring performance of these methods is limited since they do not explicitly estimate image blur. These methods are strictly local operators.
Gaussian deblurring, as one of the causes of mild blur i.e. out-of-focus is usually modeled as a Gaussian function [31]. Most of the Gaussian deblurring methods [16, 5, 58], estimate the blur kernel (i.e. Gaussian) parameters like variance, location of mean, and its orientation and then apply non-blind techniques to perform the deblurring task. However, these methods are sensitive to noise and require high inference time. Also, the error in the estimates of the Gaussian kernel parameters leads to over-sharp or over-smoothed results with artifacts [12].
A line of works [29, 41, 28, 6], achieves uniform and nonuniform deblurring tasks from the given single test input blurry image through self-supervised learning. In [41, 28], two neural networks were used to predict the sharp image and the blur kernel using a total-variation (TV) based MAP estimator. Bayesian inference implemented via Monte-Carlo sampling and [6] using deep neural network-based ensemble learning technique respectively for uniform deblurring. [29] addresses both uniform and nonuniform motion deblurring using the Monte Carlo expectation maximization algorithm. These methods require high-inference time and are computationally complex though.
Towards addressing the limitations of data dependency, time, and computational complexities for the practical use case of real-time implementation, we proposed a blind lightweight yet robust purely image data-independent model to restore the distorted image by mild blur.
@@ 3 Proposed method
In this section, the introduced RKG dataset, the proposed NSD-DIL model, dedicated Regularization, and explainable DRK are discussed in greater detail.
@@ 3.1 Problem formulation
In general, the End-to-End Deblurring (E2ED) framework can be observed as the degradation model followed by its inverse degradation model. Both the input and output of the E2ED model are sharp images. At first, the degradation model outputs a degraded version of the sharp image i.e. the blurred image, depicting the degradation process. Later, the inverse degradation model restores the sharp image from the obtained blur image i.e. the sharp image, through the restoration process. The degradation and inverse degradation models are discussed below.
Figure 2: The training and inference methodology of the proposed NSD-DIL model.
Degradation model: The degradation model typically represents the convolution operation between the input sharp image and the degradation kernel ($K$), which results in a degraded image i.e. blurred image. The mathematical representation of the correspondence between sharp image $X\in{R}^{M\times N}$ and blur image $y\in{R}^{M\times N}$ is given in Eq. (1), assuming the spatial invariant blur [27].
@@
Inverse degradation model: The objective of the deblurring task is to achieve the inverse of the degradation model i.e., to reconstruct the sharp image from the given blurred image input. The mathematical representation of the inverse degradation model is given in Eq. (2).
@@
Based on the observed E2ED model, we propose a simple, computationally efficient, and novel framework that is independent of images to learn the deblurring task. Here, we propose to learn the inverse degradation model forming an identity relation with the degradation kernel. The identity relation is given in Eq. (3). The proposed formulation simplifies the task of learning the deblurring problem from image datasets using deep architectures with several millions of learnable parameters, to the task of “identity” learning between the degradation kernel $K$ and the inverse degradation kernel $K^{-1}$ through a custom linear CNN (discussed in Sec.3.3). Hence, we term the proposed identity learning using a deep model as “Deep Identity Learning (DIL)".
@@ 3.2 Random Kernel Gallery (RKG) Dataset
The proposed NSD-DIL method is not limited to the “RKG-dataset” with anisotropic Gaussian kernels. NSD-DIL can be re-trained on any other kernel distribution if proven to be a more accurate representation of the degradations kernel space.
@@ 3.3 Linear Convolutional Neural Network (L-CNN)
The general limitations of the networks that maintain the dimensions of the input at the output, like computational time, complexity, and memory are because of operations in high dimensional space. It is important to note that these limitations are not valid in this work as the input to the CNN is kernel $K$, a matrix of very small dimension (i.e., $11\times 1$ 11), compared to very high dimensional or 3-D tensor input images, in practice. The learning and inference methodology of the NSD-DIL model is depicted in Fig. 2.
@@ 3.4 Deep Identity Learning (DIL)
The L-CNN is trained on the RKG dataset with DIL objective given in Eq. (4).
@@ 3.5 Deep Restoration Kernel (DRK)
Explainability: Black box to White box model The explicit representation of the inverse degradation kernel i.e. DRK, assists in the progress of understanding the learning mechanism of the deep model (i.e. L-CNN) for the deblurring task. The DRK, a matrix, and the L-CNN are the inverse instances of the degradation space spanned by the RKG dataset. The mesh plots of the DRK were presented in the spatial domain in Fig. 3 (b) and the top view of its magnitude spectrum in Fig. 3 (c). It can be vividly observed that the magnitude spectrum of the DRK has the characteristics of both low-pass and high-pass filters. It has a magnitude equal to one around the center (i.e. low-pass filter) and increasing magnitude away from the center (i.e. high-pass filter). The spatial domain form of DRK has a Mexican hat-like structure. Therefore, it can be noted that the extracted DRK can preserve the low frequencies and enhance the high frequencies, acting as a high-frequency-boosting filter or unsharp masking filter. Henceforth, it can be inferred that the proposed deep model can learn the actual deconvolution task, unlike the existing DL-based deblurring models which learn mapping from blurred data to sharp data distribution.
In this way, we can have a deeper understanding of the weights that were obtained from a trained deep model and explain the trained model. Therefore, the proposed NSD-DIL model for the deblurring task is no longer a black box deep learning model.
We also note that this DRK alone could also be employed for the Deblurring task as per Eq. 2 (refer Fig. 3), and the results were presented in Table 1. This makes the proposed method highly efficient for real-world applications as a single matrix with parameters in the order of 100, serves the deblurring task effectively. Also, the computational complexity was reduced from processing the blurred image through a deep model to the convolution operation between the DRK and the blurred image. The effective representation of the L-CNN facilitates the application of the regularization constraints on it easily. With DRK, it is a unique restoration process as it provides ample scope for the development of explainable models for deblurring.

### v2

@@ Abstract
In this paper, we propose to reformulate the blind image deblurring task to directly learn an inverse of the degradation model represented by a deep linear network. We introduce “Deep Identity Learning (DIL)”, a novel learning strategy that includes a dedicated regularization term based on the properties of linear systems, to exploit the identity relation between the degradation and inverse degradation models. The salient aspect of our proposed framework is it neither relies on a deblurring dataset nor a single input blurry image (e.g. Polyblur, a self-supervised method). This framework detours the typical degradation kernel estimation step involved in most of the existing blind deblurring solutions by the proposition of our Random Kernel Gallery (RKG) dataset. The proposed approach extends our previous Image Super-Resolution (ISR) work, NSSR-DIL, to the image deblurring task. In this work, we updated the regularization term in DIL based on Fourier transform properties of the identity relation, to deliver robust performance across a wide range of degradations. Besides the regularization term, we provide an explicit and compact representation of the learned deep linear network in a matrix form, called “Deep Restoration Kernel (DRK)” to perform image restoration. Our experiments show that the proposed method outperforms both traditional and deep learning based deblurring methods, with at least an order of 100 lesser computational resources. The $D^{3}$ model, both L-CNN & DRK, can be effortlessly extended to the Image Super-Resolution (ISR) task as well to restore the low-resolution images with fine details. The $D^{3}$ model and its kernel form representation (DRK) are lightweight yet robust and restore the blurry input in a fraction of a second.
@@
Convolution, Deblurring, Deconvolution, Degradation kernel, Optical blur
@@ 1 Introduction
Image deblurring is a well-established low-level vision task that aims to generate a sharp image from the given corresponding blurry input observation(s). In image formation, the blur degradation can be caused by various factors like, but not limited to, camera shake, observed relative motion between camera and target, camera defocus, camera lens aberration, light diffraction due to finite lens aperture, integration of light in the sensor [1]. Image deblurring is a severely ill-posed problem as multiple combinations of a sharp image and a blur kernel can result in the same blurry image. However, practical applications in prominent domains like medical imaging, satellite imaging, and surveillance demand a sharp version of the scene of interest for its analysis.
@@
Image deblurring solutions can be broadly categorized into non-blind and blind approaches. Non-blind approaches aim to restore clear images from blurry inputs when the degradation kernel is known [2, 3]. In contrast, blind deblurring approaches work with an unknown degradation kernel. The blind deblurring works either estimate both the blur kernel and the sharp image simultaneously [4, 5, 6, 7] or the blur kernel is estimated first and then non-blind deblurring is performed [8, 9, 10]. The traditional methods to address the image deblurring task using either non-blind or blind approaches include Maximum a-Posterior (MAP) approaches with a variety of data terms [11, 12] and sophisticated prior terms [13, 14, 15, 16], Variational inference methods [17, 18], Regularization techniques [5, 19] and Sparse representation methods [6, 20]. It is demonstrated that estimating the blur kernel is easier than jointly estimating the blur kernel and sharp image together [21]. Later, since 2014, remarkable improvement in the restored sharp image quality has been observed with the advent and continuous advancements of Deep Learning (DL) techniques. The recent deblurring solutions, including the pioneering works [22, 23], adopt sophisticated techniques like Residual blocks [24, 25, 26], Generative adversarial networks [27, 28], Recurrent Neural Networks [29, 30], Transformers [31, 32], and specialized loss functions [33, 34].
Despite vigorous research efforts, the existing deblurring algorithms still exhibit the inverse correlation between accuracy and computational efficiency. Hence, these methods are not very suitable for practical applications. Regardless of their computational efficiency the traditional methods (e.g.: [4, 35, 3, 15]) fail to handle a wide range of degradations and cast the deblurring problem into an iterative framework, which is computationally expensive. Additionally, the optimization problems of these lucid models limit their practical usage. In contrast, the State-of-the-Art (SotA) DL based methods are promising but are entirely data-driven and incur significant expenses in terms of time and computational resources for training and implementation. Furthermore, DL models grapple with generalization issues and encounter challenges when operating within memory-constrained environments due to their sizable footprint. Further, the deblurring performance of the model is challenged by the error-free assumption on the blur kernel. In practice, the estimated blur kernel is prone to error. Even the slight deviation in the estimated blur kernel leads to the ill-posedness of the problem as it results in a great mismatch of the deconvolution kernel followed by the notable artifacts in the deblurred image [36, 37]. The extensive dependency on datasets for prior information could limit the generalization ability of blind approaches.
For practical applications, few such solutions exist that deliver reliable deblurring performance while maintaining accuracy, speed, and computational efficiency [1, 38]. These methods are self-supervised and offer image-specific solutions. These methods compute the deconvolution kernel and apply it directly to the given input blurry image. However, the degradation kernel estimation step involved in [1, 38] is computationally expensive and required for every input. Besides, these methods await the blurry input to estimate the blur kernel and its deconvolution inverse.
@@
In this paper, we redefine the task of learning the deblurring model from image data to merely computing an inverse of the degradation kernel space. Here, the degradation model is represented with a wide set of anisotropic Gaussian kernels perturbed with noise called as ‘Random Kernel Gallery (RKG)’ dataset and thus detours the error-prone blur kernel estimation step. We adopt the generalized Gaussian distribution to construct the RKG dataset since many point spread function (static blur) applications are symmetric and can be modeled by such distribution [38, 1, 35]. We introduce a novel learning strategy called “Deep Identity Learning” (DIL) that exploits the identity relation between the degradation model and its convolution inverse. This method builds on our previous work “NSSR-DIL: Null Shot Image Super-Resolution Using Deep Identity Learning”, a computationally efficient Image Super-Resolution (ISR) model. Here, in this work, we made the DIL objective function robust with the updated regularization term based on the Fourier transform properties of the identity system and spatial domain properties of the convolution operator to effectively ensure the identity system properties of our inverse degradation model. We characterize the deconvolution model using a custom lightweight Linear Convolutional Neural Network (L-CNN) to train on our DIL objective using the RKG dataset. Unlike existing blind self-supervised deblur frameworks like [1, 38], the proposed deblurring framework is image-independent and doesn’t require blurry or pair of blurry and sharp image data at any stage of the deblur task learning. Furthermore, due to the non-linear and deep nature of the existing deblur works in the literature, providing the explainability of their inverse degradation operation was not feasible or trivial. Therefore, besides the modified regularization term, we provide a compact and effective representation of the proposed L-CNN, in a matrix form, called ”Deep Restoration Kernel (DRK)” to directly apply and deconvolve the blurry input. Unlike the traditional methods that estimate deconvolution kernel for a single degradation kernel for the blurry image assuming the degradation in a specific way, we estimate the deconvolution kernel (i.e. DRK) for the degradation kernel space. Here, we consider many possible ways that an image could be blurred, which allows for a more accurate and effective way to restore the image. The effective representation of the deep model L-CNN in a matrix form i.e. DRK, provides an understanding of the learning mechanism of that deep model (more details are provided in section 3.5). The proposed work can be considered as the “model-aided deep learning framework” for the deblur task unlike the current SotA works that do “distribution mapping from degraded to clean image data” using DL frameworks. In our experiments, our proposed computationally efficient L-CNN and its effective representation DRK were found to perform the deblur tasks across a wide range of degradations more effectively. The highlights of our proposition are as follows.
@@
    To the best of our knowledge, $D^{3}$ is the first DL-based image data-independent deblurring framework.
@@
    We develop the deblurring model without blurry input image and its corresponding degradation prior.
@@
    We propose a novel learning function i.e. DIL based on identity system properties with an updated regularization term driven by the spatial and frequency domain inferences to effectively learn an inverse degradation model.
  4. 4.
@@
    The proposed $D^{3}$ model and the DRK perform superior deblurring, against the existing SotA deblur methods while being at least 100 times more computationally efficient.
  6. 6.
    We do demonstrate that the proposed $D^{3}$ framework can be readily extended to the Image Super-Resolution (ISR) task.
The overall organization of the rest of the paper is as follows. Section 2 presents Related work. A detailed explanation of our proposed $D^{3}$ method is given in Section 3. In Section 4, a thorough experimental setup and results are presented, followed by a Discussion in Section 5 and a Conclusion in Section 6.
@@ 2 Related work
This section presents an overview of the existing related deblurring works.
Image sharpening methods are simple and intuitive to generate sharp images. These methods share a common framework, the input image is filtered and then a proportion of the residual image is added back to the original input. This procedure boosts high-frequencies and increases the contrast in the generated image. The most well-established works include Unsharp Masking and its adaptations. These are simple and effective but sensitive to noise and generate artifacts [39, 40, 41]. The bilateral filter [42] and its variants [43, 44] were used to enhance the local contrast and high-frequency details in an image assuming the filtered output is a local transform of the guidance image. Here, the guidance image can be the same input blurry or any other similar texture image that is not heavily degraded. Based on local structure and, sharpness information [45] proposed an adaptive sharpening method that performs simultaneous noise reduction and sharpening. However, the deblurring performance of these methods is limited since they do not explicitly estimate image blur. These methods are strictly local operators.
Gaussian deblurring, as one of the causes of out-of-focus is usually modeled as a Gaussian function [46]. Most of the Gaussian deblurring methods [47, 48, 49], estimate the blur kernel (i.e. Gaussian) parameters like variance, location of mean, and its orientation and then apply non-blind techniques to perform the deblurring task. However, these methods are sensitive to noise and require high inference time. Also, the error in the estimates of the Gaussian kernel parameters leads to over-sharp or over-smoothed results with artifacts [50].
Self-supervised learning works [37, 51, 52, 53], achieves uniform and nonuniform deblurring tasks from the given single test input blurry image. In [51, 52], two neural networks were used to predict the sharp image and the blur kernel using a total-variation (TV) based MAP estimator. Bayesian inference implemented via Monte-Carlo sampling and [53] using deep neural network-based ensemble learning technique respectively for uniform deblurring. [37] addresses both uniform and nonuniform motion deblurring using the Monte Carlo expectation maximization algorithm. These methods require high-inference time and are computationally complex though. Polyblur [1] is a blind and self-supervised approach that uses polynomial functions to restore image details blurred by mild camera shake or motion blur supporting real-time applications. It estimates the blur kernel of the image by analyzing the distribution of the gradient in the image. Then it applies the kernel multiple times, based on the severity of the blur. The work 1Shot-MaxPol [38], estimates the degradation kernel using scale-space analysis in the frequency domain of the input blurry image. Then finds the deconvolution kernel, as a linear combination of finite impulse response derivative filters, that can directly convolve with the blurry input using a closed-form solution.
Towards addressing the limitations of data dependency, time, and computational complexities for the practical use case of real-time implementation, we proposed a blind lightweight yet robust purely image data-independent model to restore the blurry input image.
@@ 3 Proposed method
In this section, the introduced RKG dataset, the proposed $D^{3}$ model, dedicated Regularization, and explainable DRK are discussed in greater detail.
@@ 3.1 Problem formulation
In general, the End-to-End Deblurring (E2ED) framework can be observed as the degradation model followed by its inverse degradation model. Both the input and output of the E2ED model are sharp images. At first, the degradation model outputs a degraded version of the sharp image i.e. the blurry image, depicting the degradation process. Later, the inverse degradation model restores the sharp image from the obtained blur image i.e. the sharp image, through the restoration process. The degradation and inverse degradation models are discussed below.
Figure 2: The training and inference methodology of the proposed $D^{3}$ model.
Degradation model: The degradation model typically represents the convolution operation between the input sharp image and the degradation kernel ($K$), which results in a degraded image i.e. blurry image. The mathematical representation of the correspondence between sharp image $X\in{R}^{M\times N}$ and blur image $y\in{R}^{M\times N}$ is given in Eq. (1), assuming the spatial invariant blur [21].
@@
Inverse degradation model: The objective of the deblurring task is to achieve the inverse of the degradation model i.e. to reconstruct the sharp image from the given blurry image input. The mathematical representation of the inverse degradation model is given in Eq. (2).
@@
Based on the observed E2ED model, we propose a simple, computationally efficient, and novel framework that is independent of images to learn the deblurring task. Here, we propose to learn the inverse degradation model forming an identity relation with the degradation kernel. The identity relation is given in Eq. (3). The proposed formulation simplifies the task of learning the deblurring problem from image datasets using deep architectures with several millions of learnable parameters, to the task of “identity” learning between the degradation kernel $K$ and the inverse degradation kernel $K^{-1}$ through a custom linear CNN (discussed in Sec.3.3). Hence, we term the proposed identity learning using a deep model as “Deep Identity Learning (DIL)”.
@@ 3.2 Random Kernel Gallery (RKG) Dataset
The proposed $D^{3}$ method is not limited to the “RKG-dataset” with anisotropic Gaussian kernels. $D^{3}$ can be re-trained on any other kernel distribution if proven to be a more accurate representation of the degradations kernel space.
@@ 3.3 Linear Convolutional Neural Network (L-CNN)
The general limitations of the networks that maintain the dimensions of the input at the output, like computational time, complexity, and memory are because of operations in high dimensional space. It is important to note that these limitations are not valid in this work as the input to the CNN is kernel $K$, a matrix of very small dimension (i.e. $11\times 1$ 11), compared to very high dimensional or 3-D tensor input images, in practice. The learning and inference methodology of the $D^{3}$ model is depicted in Fig. 2.
@@ 3.4 Deep Identity Learning (DIL)
The L-CNN is trained on the RKG dataset with the DIL objective given in Eq. (4).
@@ 3.5 Deep Restoration Kernel (DRK)
Explainability: The explicit representation of the inverse degradation kernel i.e. DRK, assists in the progress of understanding the learning mechanism of the deep model (i.e. L-CNN) for the deblurring task. The DRK, a matrix, and the L-CNN are the inverse instances of the degradation space spanned by the RKG dataset. The mesh plots of the DRK were presented in the spatial domain in Fig. 3 (b) and the top view of its magnitude spectrum in Fig. 3 (c). It can be vividly observed that the magnitude spectrum of the DRK has the characteristics of both low-pass and high-pass filters. It has a magnitude equal to one around the center (i.e. low-pass filter) and increasing magnitude away from the center (i.e. high-pass filter). The spatial domain form of DRK has a Mexican hat-like structure. Therefore, it can be noted that the extracted DRK can preserve the low frequencies and enhance the high frequencies, acting as a high-frequency-boosting filter or unsharp masking filter. Henceforth, it can be inferred that the proposed deep model can learn the actual deconvolution task, unlike the existing DL-based deblurring models which learn mapping from blurry data to sharp data distribution.
In this way, we can have a deeper understanding of the weights that were obtained from a trained deep model and explain the trained model. Therefore, the proposed $D^{3}$ model for the deblurring task is no longer a black box deep learning model.
We also note that this DRK alone could also be employed for the Deblurring task as per Eq. 2 (refer Fig. 3), and the results were presented in Table 1. This makes the proposed method highly efficient for real-world applications as a single matrix with parameters in the order of 100, serves the deblurring task effectively. Also, the computational complexity was reduced from processing the blurry image through a deep model to the convolution operation between the DRK and the blurry image. The effective representation of the L-CNN facilitates the application of the regularization constraints on it easily. With DRK, it is a unique restoration process as it provides ample scope for the development of explainable models for deblurring.